OpenStudy (anonymous):

Convert the following decimal into a fraction .0.08

OpenStudy (anonymous):

8%

OpenStudy (ilovecomputers1):

8%

OpenStudy (anonymous):

COPIER @ILoveComputers1

OpenStudy (anonymous):

fraction sorry

OpenStudy (ilovecomputers1):

8/100, reduces to 2/25

OpenStudy (anonymous):

Multiply and divide 0.08 by 100 to obtain percentage. 0.08 = 8 / 100 = 8% Reduce the fraction 8 / 100 dividing numerator and denominator by 4 to obtain a reduced fraction. 8 / 100 = 2 / 25